OpenStudy (anonymous):

What is ¾ x 1 ½ ? [a] 1 ⅛ [b] 1 ¼ [c] 1 ½ [d] 1 ¾

OpenStudy (anonymous):

3/4*3/2=9/8 a

OpenStudy (anonymous):

a

OpenStudy (anonymous):

ty :) 1 more pls

OpenStudy (anonymous):

What is 4 ½ x 1 ¾ ? [a] 7 ⅛ [b] 7 ⅜ [c] 7 ⅝ [d] 7 ⅞

OpenStudy (anonymous):

9/2*7/4=63/8

OpenStudy (anonymous):

D

OpenStudy (mathmale):

Alex: Guide the other person towards his / her finding his / her own solutions and choosing the correct alternative from among the answers. Giving out answers doesn't help much in the long run. Thank you.